OpenStudy (anonymous):

In a galvanometer there is a deflection of 10V divisions per mA. The internal resistance of a galvanometer is 50ohm. If a shunt of 2.5ohm is connected to the galvanometer, calculate the maximum current which the galvanometer can read.

OpenStudy (kropot72):

Let resistance of the parallel combination of the galvanometer and the shunt = R(p) \[R(p)=\frac{2.5\times 50}{2.5+50}\] The maximum current the galvanometer can read with the shunt connected is: (maximum current without shunt) * 50/R(p)

OpenStudy (anonymous):

sorry brother but the right answer is 125mA

OpenStudy (kropot72):

I did not give an answer. This part of the question does not make sense "there is a deflection of 10V divisions per mA". You will note I did not give a value for the maximum current without shunt.

OpenStudy (anonymous):

ya i think you are right but this question was given by my sir in exam and he made no correction to it...

OpenStudy (kropot72):

The current multiplying factor 50/R(p) is correct using the resistance values in the question.

OpenStudy (vincent-lyon.fr):

This should read: "In a galvanometer there is a deflection of 10 divisions per mA" What is the maximum reading? 100 divisions?

OpenStudy (anonymous):

I am afraid that without the max deflection in mA (or Volts), there is no way to solve it

OpenStudy (kropot72):

According to the questioner the maximum current to be read with the shunt connected is 125 mA. Therefore the maximum current to be read without the shunt would be \[\frac{125\times 2.381}{50}=5.95mA\]